ONY I TUWU TU TU 000 L alu off? 23. Calculating EAR (L04) Bergheims Quick Loans Inc. offers you three for fi knock on your
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

4=3*(1+r/100)

(1+r/100)=(4/3)

r=1.3333-1

=33.33%(Approx).

APR=33.33%*52 weeks

=1733.33%(Approx).

EAR=[(1+APR/m)^m]-1
where m=compounding periods

=[(1+0.3333)^52]-1

=313,916,515.70%(Approx).